### v4

# Counterfactual Effect Generalization: A Combinatorial Definition

Abstract
The widely used ’Counterfactual’ definition of Causal Effects was derived for unbiasedness and accuracy - and not generalizability. We propose a Combinatorial definition for the External Validity (EV) of intervention effects. We first define the concept of an effect observation ’background’. We then formulate conditions for effect generalization based on their sets of (observable and unobservable) backgrounds. This reveals two limits for effect generalization: (1) when effects are observed under all their enumerable backgrounds, or, (2) when backgrounds have become sufficiently randomized. We use the resulting combinatorial framework to re-examine several issues in the original counterfactual formulation: out-of-sample validity, concurrent estimation of multiple effects, bias-variance tradeoffs, statistical power, and connections to current predictive and explaining techniques.
Methodologically, the definitions also allow us to also replace the parametric estimation problems that followed the counterfactual definition by combinatorial enumeration and randomization problems in non-experimental samples. We use this non-parametric framework to demonstrate (External Validity, Unconfoundness and Precision) tradeoffs in the performance of popular supervised, explaining, and causal-effect estimators. We demonstrate the approach also allows for the use of these methods in non-i.i.d. samples. The COVID19 pandemic highlighted the need for learning solutions to provide predictions in severally incomplete samples. We demonstrate applications in this pressing problem.
1 Introduction
Suppose a sample has a Data Generating Process (DGP) which we can only partially observe. We would like to know under what conditions effects (e.g., of an agent’s action or an administered treatment) observed in this sample can be taken as true representations of the effect of its variables. The main challenge is that of confounding of effect estimates. Confounding refers to sampling conditions where uncontrolled variation of out-of-sample or in-sample factors bias in-sample effect estimates. Under these conditions, sample effect observations (i.e., difference in outcomes among sample units) reflect not only the true effect of sample factors, but also of their uncontrolled and unobserved factors. Under a broad class of sampling conditions, effects become, in fact, fully unidentifiable. The ability of estimates to generalize across out-of-sample conditions is often described as a desirable property, but it is not always clear what is meant, and optimized for, in these cases. We study generalization, in particular, from the perspective of Experimental Design, and in scenarios with time-varying treatment exposure and confounding. Formulation of specific sampling conditions is the starting point of many theories of causation (Sect.1.1 Effect estimates and Unobserved Variation), but are often abandoned in lieu of complex regression and parametric methods. This last step makes it difficult to formulate crucial statistical properties for effect estimates such as their precision, biases and out-of-sample generalizability. These requirements have been barely established for popular counterfactual-based effect estimation methods, like Shapley value black-box regressions  [Burkart and Huber, 2021, Lundberg and Lee, 2017, Aas et al., 2021] and g-formula  [Chatton et al., 2020], or most supervised prediction methods. In practice, this makes them difficult to use, and for users to determine when their outputs can be trusted (does the sample have enough observations? variables?). We provide definitions and demonstrate how to select subsamples with simultaneous high generalizability and low confoundness. We demonstrate uses when estimating causal effects non-parametrically, or understanding the performance of out-of-sample supervised predictors and black-box explainers. The latter is currently an intensely studied subject in Artificial Intelligence and Machine Learning  [Lundberg and Lee, 2017, Burkart and Huber, 2021]. This article continues progress on the challenge of estimating causal effects non-parametrically, exclusively from effect observations  [F. Ribeiro et al., 2022].
Partition the set $Z$ of DGP variables into two subsets, $Z=X_{m}\cup U_{q}$, the first with $m$ observed variables $X_{m}=\{-1,+1\}^{m}$, and the second with $q$ unobserved $U_{q}=\{-1,+1\}^{q}$. Each sample unit is, consequently, characterized by a subset of observed factors, $x\in X_{m}$. Imagine also the DGP is constrained by precedence relationships among its variables, described by two partially ordered set (posets) $G(Z)=\{G(X),G(U)\}$. The distinct ways in which $\{G(X),G(U)\}$ can be combined is described by the distinct linear extensions among these posets. For poset chains $\{G(X),G(U)\}=\{\;\mathbf{a}<\mathbf{b},\;\;u_{1}<u_{2}<u_{3}\;\}$, for example, the extensions are
Example 1 ($U{-}X$ Linear Extensions).
$\mathbf{ab}u_{1}u_{2}u_{3}\,(*),\,\mathbf{a}u_{1}\mathbf{b}u_{2}u_{3},\allowbreak\,\mathbf{a}u_{1}u_{2}\mathbf{b}u_{3},\allowbreak\,\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b},\allowbreak u_{1}\mathbf{ab}u_{2}u_{3}\,(*),\,u_{1}\mathbf{a}u_{2}\mathbf{b}u_{3},\,u_{1}\mathbf{a}u_{2}u_{3}\mathbf{b},\,u_{1}u_{2}\mathbf{ab}u_{3}\,(*),\,u_{1}u_{2}\mathbf{a}u_{3}\mathbf{b},\,\allowbreak u_{1}u_{2}u_{3}\mathbf{ab}\,(*),\,\allowbreak\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b}.$
The different extensions can be seen as different confounding conditions under which sample $X$ can be observed. We imagine real-world samples can be collected under any such conditions, and we would like to estimate the statistical properties of effects across partially observed samples, $U\neq\emptyset$. The likelihood of observation under a $U{-}X$ extension follow, in turn, relationships between the distributions of factors in $X$ and $U$,
Example 2 ($U{-}X$ Data-Generating Order Relations).
[…]
where $p_{t}(b\,|\,a)$ and $p_{t}(u_{i}\,|\,a)$ denote the probability of changes in the value of binary factors $a\in X$ and $u_{i}\subseteq U$, given the observation of a $a$ at a time $t=0$, and $\leq$ denotes poset comparability. The first condition implies that the partial order $a<b<Z-\{a\}$ will apply with high likelihood in the resulting sample. The condition correspond to the first extension in Example.1. That is, that there are no confounders between subsequent $a$ changes, and observed effects in the sample emulate, from frequency-based constraints, a typical counterfactual estimator (Sect.1.1 Effect estimates and Unobserved Variation). This is, however, only true for specific subsamples in the other two scenarios (in equal proportion across all factors in the second and in rare subsamples in the third). For short, we refer to $U{-}X$ systems described by the first conditions as under fast sampling, and the last as under slow sampling. We define alternative sample conditions for the (1) separability (whether pairs of causes and confounders can be distinguished) using only effect differences (or their variance) in samples, and more restrictive conditions of (2) identifiability of causes. Of particular interest in real-world large-scale data is the condition where $U{-}X$ relationships (’treatment-background’ assignments) stay constant, while (one of these) subsets are allowed to change freely. Under some extensions, treatment-background assignment in $X$ can become random-like, and independent of unobserved confounders. These considerations lead to a theoretical model to describe how combinatorial and ordinal relations in $U{-}X$ systems impact the generalizability and confoundness in samples’ effect observations and the performance of estimators that use them.
1.0.1 Alternative Orders and Confounding: A Preliminary illustration
Let’s consider the observed effects of sample factors that are either causes or confounders in a sample with a true cause $a$, a confounder $b$ (with effects $y$ completely correlated with its root cause $a$), and a spurious variable $c$, $X=\{a,b,c\}$. Let’s consider, in particular, differences in observed effects of $a$ and $b$ across different values of other factors. We will use Latin square diagrams as aid to illustrate several concepts in this article. Fig.1(a) shows an example with $4$ factors. Each square is associated with a sample unit $x_{0}$, its reference unit, which stratifies a population (other units) across its cells. A square with reference $x_{0}$ correspond to the set of effect observations with singleton factor differences,
$\begin{split}\bigcup_{x\in\mathcal{P}(X)}{{\Delta}}y(x-x_{0})=\{a\},\qquad\forall\,a\in X.\end{split}$
where $\mathcal{P}(X)$ is the powerset (set of sets) of $X$ and ${\Delta}y(x-x_{0})$ is an effect observation in outcome-of-interest, $y\in\mathbb{R}$, after an observed intervention on factors $x-x_{0}$. A single square contains therefore $\binom{m}{2}$ effect observations, and a sample (or all its squares) is limited to $2^{m}$ observations. We will say that the square’s diagonals enumerate, for example, observations of effect for a fixed factor across different (observed) backgrounds. The diagonal effect observations for the cause $a$ and confounder $b$ are
$\hfil\textrm{diagonal}\;a\hfil&\hfil\textrm{diagonal}\;b\hfil\\[8.0pt] \Delta{y}(a)&&\hfil\hskip 30.&\Delta{y}(b)&&\\ &\Delta{y}(a\,|\,b){-}\Delta{y}(a){=}0&\hfil\hskip 30.&&\Delta{y}(b\,|\,a){-}\Delta{y}(b){=}\Delta{y}(a)&\\ \Delta{y}(a\,|\,b)&&{=}0,\hfil\hskip 30.&\Delta{y}(b\,|\,a)&&{=}{-}\Delta{y}(a),\\ &\Delta{y}(a\,|\,bc){-}\Delta{y}(a|b){=}0&\hfil\hskip 30.&&\Delta{y}(b\,|\,ac){-}\Delta{y}(b\,|\,a){=}0&\\ \Delta{y}(a\,|\,bc)&&\hfil\hskip 30.&\Delta{y}(b\,|\,ac)&&\\ \end{array}$
We see that when the cause $a$ is in the sample, it sustains its effects through arbitrary sample imputations (indicated by the conditional statements), while the confounder also assumes negative values. In a square and fully varying set $X$ of factors, confounders change their effect observations several times, while causes remain invariant throughout all enumerated variation.
More importantly, however, this indicates that, to identify confounders in fully-observed samples ($U=\emptyset$), we need samples where we, minimally, observe the effect of both adding factor $a$ to units with $b$, and of adding factor $b$ to units with $a$ (i.e., under alternative orders). These conditions are necessary because $b$ has no effects only in the latter case, $\Delta y(a)+\Delta y(a\,|\,b)>\Delta y(b)+\Delta y(b\,|\,a)$. Samples that include these conditions allows us to see, from effect observations alone, that $b$’s effects were brought by $b$’s correlation with $a$, and not its own independent effect on $y$. This constitutes a simple sample requirement for cases with $m=2$ and $q=0$. The rest of this article generalizes this to the $m$ factor case and the possibility of $q$ unobserved factors.
Principle 1 (Cause-Confounder Separability).
[…]
According to the principle, under some ordinal conditions on $U{-}X$ factors, effect observations for confounders have an expected effect variance of $0.25\times\rho_{ab}$ (illustrated by the number of times the confounder appears before its root cause in square rows) and causal variables 00 of the true effect $\Delta y(a)$. One such condition is of a fully observed sample, $U=\emptyset$, when the principle allow the non-parametric identification of causes. In this case, the principle is associated with the observation of each effect under the cyclic permutation of all others factors (as illustrated by square rows), and it says that effect variances, ${\rm I\kern-1.99997ptE}[\,\textrm{Var}[\Delta\hat{y}(a)]\,]$, are a sufficient statistic, for a fixed level of cause-confounder correlation. Other conditions, established below, are required otherwise. These conditions correspond, non-parametrically, to the observation of multiple squares.
1.1 Effect estimates and Unobserved Variation
Before formulating necessary conditions for the generalizability and confoudness of effect observation in samples, we give our working definitions of effects and their generalizability. Subsequent sections justify, with increasing mathematical rigor, these combinatorial definitions. According to the Potential Outcomes framework’ definitions for what is a causal effect  [Morgan and Winship, 2007, Rubin, 2005], if $y$ is an outcome of interest and $a$ is a treatment indicator, then the causal effect of $a$ is the difference
$\Delta y(a)=y_{i}^{\texttt{+}a}-y_{i}^{\texttt{-}a},$ (1)
where $y_{i}^{\texttt{+}a}$ is the outcome of individual $i$ under the treatment, and, $y_{i}^{\texttt{-}a}$ without the treatment. The central concept behind Eq. (1) was inspired by experimental estimation: by fixing every factor, other than the treatment, we can declare that the observed difference in outcome was certainly caused by the treatment, and the treatment alone. The definition is an ideal, as it is impossible to observe outcomes for an individual, concurrently, in two different and totally fixed conditions. The theory goes that we may, instead, ’fix’ factors in expectation, and across individuals. If the treated and non-treated subpopulations have the same expected values across all factors then any difference between the groups is due to the treatment, given large enough samples. This can be paraphrased with an independence statement: treatment assignment must be independent on all outcome-relevant factors. This rationale led to the notion of Sample Balance in non-experimental estimation  [Rubin, 2005, Morgan and Winship, 2007] and the objective most current counterfactual causal effect estimators maximize.
While not a lot is known theoretically about the Counterfactual definition  [Abadie and Imbens, 2006, Shalit et al., 2017], we can observe that for Eq.(1) and a pair of sample units $x_{i}$ and $x_{j}$,
$\textrm{Var}(x_{i}{-}x_{j})=\textrm{Var}(x_{i})+\textrm{Var}(x_{j})-2\textrm{Cov}(x_{i},x_{j}).$ (2)
A maximally accurate estimator (i.e., one with minimum variance) minimizes the covariance between sample units it is comparing. There is, however, a problem with this definition. It is only guaranteed to hold under a very strict set of conditions. Namely, it holds only if all factors of relevance are held constant among individuals. Such estimates are poised to not generalize across populations.
We consider an effect, instead, as
$\Delta y(a)$ $=\frac{1}{m}\sum\limits_{\pi\in\Pi_{n}(Z/\{a\})}\Big{[}\,y_{>}(a\,|\,\pi)-y_{\leq}(a\,|\,\pi)\,\Big{]},$ (3)
where $y_{\leq}$ is the observed outcome $y$ of a sample units with, exclusively, the set of factors before $a$ in the permutation order $\pi$, and $y_{>}$ with factors after $a$. Furthermore, $\Pi_{n}(Z/\{a\})$ is a subset of the full set of permutations of factors $Z$ observable between sequential changes in $a$ values, and is determined by the cycle-structure of permutations over $Z/\{a\}$. Sample ordinal patterns, correlations and periodicity determine $\Pi_{n}$. Difference between units in Eq.(3) constitutes observation of effects of the same factor in distinct backgrounds $\pi\in\Pi_{n}(Z/\{a\})$. Each background fixes a set of factors for their ’treatment’ (before $a$ in $\pi$ or, not containing $a$) and ’control’ (after $a$) for sample units. For a fixed order $\pi$, there are $m$ orders that preserve the relative order of in-sample factors. The rows of a Latin-Square diagram, Fig.1(a), describe a single such order for factors $X$. To see this, choose any factor in a square diagram and follow its subsequent factors cyclically (e.g., $b<c<d$ for $a$). The set of enumerable permutations will be generalized to samples, $\Pi_{n}(U=Z/X)$, in Sect.3 Sampling Treatment-Background Assignments.
At the limit, when all factors in $U=Z/X$ are independent, $\Pi_{n}$ correspond to the full set of permutations of size, $|\Pi_{n}|=q!$. At the same time, when all factors are serially correlated, $\Pi_{n}$ correspond to to only one, $|\Pi_{n}|=1$. Sample correlations and fixations in the cycle structure of permutations can drastically reduce $|\Pi_{n}|$, many times asymptotically and even in very large samples. In the next section, we will use sample ordinal relations to describe this relationship in further detail. We will consider how incompleteness of $\Pi_{n}$ in respect to both $X$ and $U$ affects (often in unknown ways) the generalizability of effect estimates. Based on these conditions, we say an effect is Externally Valid (EV) if it holds under the system $U{-}X$ set of unobserved backgrounds,
$\textrm{EV}(a)=\textrm{Var}^{-1}\Big{[}\Delta y(a)\,|\,\Pi_{n}(Z/\{a\})\Big{]}.$ (4)
EV is here the inverse of the variance (i.e., the precision) of observed effects under all asymptotically observable backgrounds, $\Pi_{n}(Z/\{a\})$. The most important element of these definitions are $\Pi_{n}(Z/\{a\})$, subsets of unobservable permutations in a $\{X,U\}$ when $a$ is fixed. This defines causal effects in a way that is almost opposite to Eq. (1). It calls for effects to be observed under large variation, as opposed to no variation. According to Eq.(3,4), the estimate is, in fact, the one with minimal EV. Underlying these definitions and Principle.1 is the observation that, unlike confounders, causes can be defined fundamentally as ’perfect controls’ [Pearl, 2000]. Similar definitions have been explored in work based on robustness (of models) [Peters et al., 2016, Buehlmann, 2020] (Sect.4 Related Work). The non-parametric definitions here frame the problem, in particular, not as simply the invariance (of effects), but invariance of effects under maximum variance of their ’backgrounds’. Shifting the focus to the later, $\Pi_{n}(U{=}X/Z)$, is a key but challenging part of the proposed solution.
Figure 1: (a) $4{\times}4$ Latin-Square (’square’) as sets of effect observation for sample unit $x_{0}\subseteq X$, (b) $3$ samples with increasing random factors $X$ and their ordinal statistics (interval ticks), $m=\{1,2,4\}$, and alternative sampling of confounding factors $U$ (circles)(top), resulting distribution and percentage of enumerable permutations of effect observations across ordinal variables (bottom and right).
2 Confounding Ordinal Statistics and their Unobserved Gaps
If $\{X_{a},X_{b},\ldots,X_{[m]}\}$ is a random $m$-sample from a particular distribution, it will be important to know whether a factor $b$ follow or precede sequential changes in $a$ (e.g., events $a_{1},a_{2}$). For example, whether $a_{1}<b<a_{2}$ or $a_{1}<a_{2}<b$ lead to distinct scenarios in the estimation of $a$’s effect. The likelihood of such orders is determined by the DGP, as illustrated in Example.2. To that end, we define $X_{(1)},X_{(2)},\ldots,X_{(m)}$ to be a sorted version of the sample  [Gut, 2009], that is,
$X_{\min}\equiv X_{(1)}<X_{(2)}<\ldots\leq X_{(m)}\equiv X_{\max}$
Notice that $X_{\min}$ is a random variable. For simplicity, we assume that, in sufficiently large samples and time resolution, the order values cannot coincide, leading to strict inequalities (as opposed to $\leq$).
Suppose that $X_{1},X_{2},\ldots,X_{m}$ follow a continuous distribution with PDF $p$ and CDF $F$. We will use $X$, $U$ or $Z$ as alternative labels when considering distributions over observed, unobserved or both factors. Imagine a random sample of $U$ factors falling in such a way that its minimum is below a fixed observed value $x$. Fig.1(b) illustrates examples of draws where circles are samples of $U$ and interval ticks of $X$. The intervals show $X_{(1)}$ in samples with increasing dimensionality $m$. Circles to the left of the dashed line represent irreducible sample confounding effects. Under i.i.d. sampling, this can be expressed as
$P(x\,\text{is 'confounded' by }U)$ $=P(\text{at least one factor in }U\text{ is}\leq x)$ (5)
$=P(U_{(1)}\leq x)=F_{X_{(1)}},$
which uses the observation that $x$ is the minimum if and only if all the $U$ are greater than $x$. The corresponding pdf is then $p_{X_{(1)}(x)}=m[1-F(x)]^{m-1}p(x)$.
In particular, we are interested in the distribution of the $k$-th sample gap and its associated variables,
$\Delta_{(k)}X\stackrel{{\scriptstyle\text{def}}}{{=}}X_{(k)}-X_{(k+1)}$ $,\qquad\Delta_{(k)}y(X{=}x)\stackrel{{\scriptstyle\text{def}}}{{=}}Y_{(k)}(X{=}x)-Y_{(k+1)}(X{=}x),$
$\Delta U_{(k)}\stackrel{{\scriptstyle\text{def}}}{{=}}$ $P(u_{(k)}\,|\,X_{(k)}<u<X_{(k+1)}),$ (6)
[…]
$\Delta_{(k)}m\stackrel{{\scriptstyle\text{def}}}{{=}}|\,a\in\text{{Range}}(X_{(k)})\mid P(X_{(k)}{=}a)>0\,|,$
$\Delta_{(k)}q\stackrel{{\scriptstyle\text{def}}}{{=}}|\,u\in\text{{Range}}(\Delta U_{(k)})\mid P(U_{(k)}=u)>0\,|,$
where $\Delta_{(k)}m$ and $\Delta_{(k)}q$ are the number of observed and unobserved factors with positive probability in, respectively, an ordinal position or gap. These gap statistics are closely associated with the possible extensions among posets $\{G(X),G(Y)\}$ (such as those illustrated in Example.1).
An effect statistic for a gap will become relevant because it reflects only the effect of new causes introduced within the gap ($\Delta y(u)$, $a<u_{k}<b$ for two sample variables $a$ and $b$ and unobserved causes $u_{k}\subseteq U$). Of particular interest are gap effects when in-sample factors are fixed across ordinal observation sequences. Observed outcome differences in a gap $(k)$ reflect then exclusively the effect of unobserved factors falling in $(k)$. In a square diagram with $X=\{a,b,c\}$, for example, $Y_{(k)}(X{=}a)-Y_{(k+1)}(X{=}a)$ denotes outcome differences across its main diagonal, Fig.1(a). For the case of an increasing $m$, where $U_{(1)}\rightarrow\emptyset$, Eq.(5) suggest distinct levels of confounding across sample effect observations,
Proposition 1 (Partial-Observability and Effect Variance).
[…]
This suggests that even when a factor $b$ is not observed at a minimal order $(k=1)$, the unobserved confounding effects at $(k=1)$ can still be discounted by using other sample factor effects prior to $b$ in the order. For example, when $a$ is observed at $(k=1)$ but $b$ only at $(k=2)$, $\textrm{Var}[\Delta y_{(2)}(b\,|\,a)-\Delta y_{(1)}(a)]\leq\textrm{Var}[\Delta y_{(2)}(b\,|\,a)]$.
We will consider, in particular, the relationship between distributions of causes across gaps and (1) the number of permutations that $U{-}X$ systems can generate, and (2) their impact in the statistics of effect observations across sample ordinal positions. According to the separability principle (Principle.1), the first is a requirement for the separability of $U{-}X$ causes and confounders using effect observations. Some distributions of unobserved causes across gaps can asymptotically generate all distinct orders (all observable permutations of unobserved factors) and some can’t. This can be described by a multinomial coefficient, and by comparing the cycle structure of $U{-}X$ permutations to their ideal full set of permutations $q!$,
[…]
=($={q\choose\Delta q_{(1)},\Delta q_{(2)},\ldots,\Delta q_{(m)}}^{-1}.$ (7)
We will study, in particular, $U{-}X$ systems where $\Delta_{(k)}m\gg\Delta_{(k)}q$ (fast) or $\Delta_{(k)}m\ll\Delta_{(k)}q$ (slow) in specific sampling periods. Fig.1(b, bottom-right) shows examples of this relationship in the former. The lower panel shows the limiting cases of $(\Delta m_{(1)},\Delta m_{(2)},...)=(m,m,...,)$ that can enumerate all orders (ratio $\nicefrac{{4!}}{{4!}}=1$), the case of (1,1,,)(1,1,...,)( 1 , 1 , … , ) that can enumerate only one ($\nicefrac{{1!}}{{4!}}=\nicefrac{{1}}{{4!}}$) and an intermediary case. The figure also shows the corresponding squares, and thus effect observations, available across the resulting samples.
We will demonstrate that effect generalizability assumes a particularly simple form with the following two sample gap conditions across time $t$,
$P(x<u\,|\,\Delta U_{(0)}(t)=u_{0},\Delta U_{(1)}(t)=u_{1},...,\Delta U_{(m)}(t)=u_{m})$ (8)
$=P(x<u\,|\,\Delta U_{(0)}(t+m)=u_{0},\Delta U_{(1)}(t+m)=u_{1},...)$ (fixed background)
$=P(x<u\,|\,\Delta U_{(0)}(t+m)=u_{\delta 0},\Delta U_{(1)}(t+m)=u_{\delta 1},...),$
where $u_{0}\perp u_{\delta 0}$, $u_{1}\perp u_{\delta 1}$, etc. The first corresponds to the case of fixed (unobserved) effect backgrounds for a time of at least $m$, and it can lead to generalizeable effects in fast $U{-}X$ systems. The second correspond to cases where effects are observed at time delays $\delta$ (and assignment of unobserved confounders across gaps), $\pi_{\delta}$, no longer depends on the initial background, $\pi_{0}$ (Sect. 4 Related Work). It is useful in slow systems. Crucially, to formulate the $\delta$ required for background randomization, we will consider to what extent gap statistics allows us to identify linear extensions among posets $\{G(X),G(U)\}$, which, in turn, allow us to formulate $\delta$ theoretically, given the specific extension.
Confounding $U$ in samples following the conditions in Eq.(8) can be expressed in the form
$U$ $=\frac{(q-1)}{q}\Delta U_{(1)}+\frac{(q-2)}{q-1}\Delta U_{(2)}+\ldots+\frac{(q-k)}{q-k+1}\Delta U_{(k)},$ (9)
or, similarly, by saying that the variables $U_{(k)}$ form an (additive) Markov chain. The observation that order statistics under the previous conditions form a Markov chain was first made by Kolmogorov  [Rényi, 1953]. If $U$ follow both exponential and uniform distributions gap effects will have this additive property. While this not always the case in generic $U{-}X$ non-experimental samples, decomposition of samples in subsets of effect observations with these properties will offer comparable advantages. Eq.(9) suggests that confoudness of effects in squares increases linearly across their columns, and is trivially associated with an ANOVA variance decomposition for effect variance in samples (Appendix G).
Figure 2: (a) sampling sequences across sample, $X=\{a,b,c\}$, gaps and confounder partition $U=\{\{u_{1}\},\{u_{2},u3\},\{u_{4}\}\}$, (d) normalized variation separation distance (distance to a Uniform distribution) of all backgrounds $\pi_{0}$ to other backgrounds after time $t$, vertical dashed lines mark limits on successive background randomness based on their inversions (top panel shows the same for hamming distances across all sample effect observations), (c) randomization in graphical models of in-sample treatment $A$, other factors $W$, unobserved factors $U$, and outcome $Y$, (d) combinatorial relations in a square ($m=2$), circles are differences and gray circles intersections of indicated letters between $x_{0}$ and other units.
3 Sampling Treatment-Background Assignments
We will start with a definition for generalizeable sample effect sequences, as permutations of remainders of $Z/X=U$ for a fixed $X$ with unobserved factors $U$. This will formulate what sampling sequences, and thus subsets of permutations $\Pi_{n}(Z/X)$, are necessary to make claims about the generalizability of effects in $X$. Eq.(3) formulated effects for a single fixed factor $a$, where backgrounds were $\Pi_{n}(Z/\{a\})$, and is generalized here for samples $X$. The next results rely on the following assumptions, which are somewhat typical,
Assumptions 1.
  1. 1.
    Constant DGP. The set of observed, $X$, and unobserved factors, $U$, in $Z=X\cup U$ do not change during the sampling period. This implies no measurement errors, and that, conditioned on all other factors in $Z$, probabilities and effect of causes do not change or have low variance. This leads to typical notions of irreducible variance and errors in samples  [Tibshirani et al., 2001].
  2. 2.
    Multiple Exposures. Effects remain constant under multiple exposures. While this is implied by most definitions of effects, it is inadequate for continuous and cumulative effect interventions, such as treatments involving dosages and other interactions with the numbers of exposures.
We also make two simplifying assumptions,
Assumptions 2.
  1. 1.
    Large-scale data. The sample $X$ includes both causes and spurious factors. In our cases of interest, such large-scale economic, genetic and medical data these are typically available.
  2. 2.
    Non-strongly antagonistic interactions. Principle.(1) imply additive effects among causes (see Appendix.B for a proof). This is a weaker version of an assumption of additivity of unobserved causes’ effects, useful in samples that do contain causes in such ideal conditions [^1] Many times, true causes are unobserved but we still want to find the closest surrogate in-sample factor that reproduces the cause effects across observable conditions. In these cases, additivity or monotonicity of effects might be too strong a requirement, and other non-interaction requirements for unobserved causes and gaps can be formulated. These requirements are unnecessary for Principle.1, except for strongly antagonistic effects, and unnecessary in the empirical cases we have considered so far. Interaction effects for in-sample effects are adjusted by averaging across square in-sample backgrounds, Eq.(3). .
Imagine our effect observations to be discrete and equally spaced points in time. If an effect sample at time $t$ is next sampled at time $t^{\prime}$, we write $b(t)=t^{\prime}$. If $t$ is an integer then let $t\bmod q$ denote the remainder when $t$ is divided by $q$. Finally we let $\bar{q}$ be a shorthand for the set of possible remainders $\{0,1,2,...,q-1\}$. Sampling sequences can be interpreted mathematically in several ways. We choose to think of them as permutations $\pi:b(t)$ of the set $K$ of integer sample gap indices. We call each factor $k$ in $\bar{q}$ an individual sample gap index, $u_{k}$, in reference to Eq.(2), and we call a permutation of $\bar{q}$ a single background. In respect to sample effects, each permutation defines, in turn, a distinct treatment-background assignment. Regularities in this assignment in $U{-}X$ systems defines their treatment-background mechanisms. Thus a sampling strategy defines a function $b:K\rightarrow K$, where $b(t)$ is the next sample gap to be observed at time $t\in\mathbb{N^{+}}$. Remember permutations can have fixation points. Each such effect observation correspond to a subset of unobserved factors $u_{k}\subseteq U$, Eq.(2), and we say the effect of factor $a$ is observed in the background of $u_{k}$ in this case.
Definition 1 (Sampling Sequence).
A sampling sequence over $X$ is a finite sequence $b$ of nonnegative integers $b(0),b(1),...,b(q-1)$ such that the mapping $b(t):K\rightarrow K$ defined by $b(t)=t+b([t]_{q})$ is a permutation of out-sample factors (’background’ factors). A sampling sequence thus determines, for each time $t$, a surjective $U{-}X$ background-treatment assignment.
Fig.2(a) shows a sample $X_{3}=\{a,b,c\}$ with a fixed order statistic, $a<b<c$, and changing $U_{4}=\{u_{1},u_{2},u_{3},u_{4}\}$ (slow sampling). The diagram shows the transition of backgrounds under repeated observations of effects with a fixed-order $a<b<c$. Backgrounds are partitioned in this example as $\{\{u_{1}\},\{u_{2},u_{3}\},u_{4}\}$. Each arrow correspond to the sequential mapping of these unobserved factors (arrow labels) across sample gaps (intervals).
Effect backgrounds $Z/X$ are not necessarily constant due to the distinct frequencies, correlations and periodicities among $U$ and $X$ factors. A key definition in this article is
Definition 2 (EV-increasing sequence).
A finite sampling sequence $b$ of nonnegative integers $b(0)...b(q-1)$, and its $U{-}X$ assignments, is EV-increasing if and only if the integers $[t+b(t)]_{q}$ are distinct.
In Fig.2(a), for example, the top and bottom diagrams illustrate sequences that are not EV increasing, as each variable effect is observed in only one background. The middle diagram is EV increasing with period of $2$, as effects are observed in new backgrounds at each time, and all enumerable backgrounds for a fixed $Z/X$ are observed after time $2$. A complementary property for sampling sequences is
Definition 3 (CF non-increasing sequence).
A finite sampling sequence $b$ of nonnegative integers $b(0)...b(q-1)$, and its $U{-}X$ assignments, is confounding bias non-increasing, or CF non-increasing, if and only if the backgrounds $[t+b(t)]_{q}$ are the same across all sample factors $X$.
In Fig.2(a), the top and middle diagrams are CF non-increasing. The bottom is CF non-increasing for a period of $2$. The ***-marked extensions in Example.1 are CF non-increasing, but so are any sequences with cyclic permutations of in-sample factors under the condition that the unobserved background is fixed (i.e., subsequent rows of sample squares).
Definition 4 (EV-CF sequences).
A finite sampling sequence $b$ of nonnegative integers $b(0),...,b(q-1)$ is EV-CF if and only if it is both EV increasing and CF non-increasing. We say sample effect observations are ’generalizeable’ under EV-increasing sequences, and ’unconfounded’ in EV-CF sequences.
This leads to the notion of enumeration of effect backgrounds,
Definition 5 ($U{-}X$ Effect Background Enumeration).
[…]
That’s because, for a fixed gap index $k$, it takes $2^{t}\times k(\bmod 2n-1)=k$ time for $u_{k}$ to get back to its original position in this case [Diaconis et al., 1983], and for all background factors to be enumerated222as result of the recursion $(2(2...(2k(\bmod 2n-1))...)))$ for each gap position $k$.. This can be restated as
$2^{t}(\bmod 2n-1)=2^{t}\ (\bmod\ q-1)=1,$ (10)
for a initial position $k=1$.
The impact of the observation of increasing backgrounds $Z/X$ on sample effects (and background enumeration) depends on the specific extension between posets $G=\{G(X),G(U)\}$, which we exemplified in Example.1, described as the partition of $U$ induced by sample gaps, Eq.(2), and now describe stochastically with a multinomial distribution, $p=\{p_{1},p_{2},...,p_{m}\}$. Each distributions (and the linear extensions they imply) is associated with a number of permutations of $U$ they can generate under a fixed $X$. This was formulated in Eq.(2), but can be made more precise by fixing an initial permutation $\pi_{0}$ and counting the number of inversions, or pairs of elements that are in the reverse order, from $\pi_{0}$. The expected number of permutation inversions, $I(\pi_{0},t)$, generated by a fixed distribution $p$ and initial background $\pi_{0}$ (Appendix.A) after time $t$ is
${\rm I\kern-1.99997ptE}[\,I(t\,|\,p,\pi_{0})\,]$ $=\frac{1}{2}\times\binom{m}{2}\times\Big{(}1-\Big{[}p_{1}^{2}+p_{2}^{2}+...+p_{m}^{2}\Big{]}^{t}\Big{)}$ (11)
where $\binom{m}{2}$ denotes the number of effect observations in a single square. For the case of $p=\{0.5,0.5\}$ the equation reduces to
${\rm I\kern-1.99997ptE}[\,I(t\,|\,\pi_{0})\,]$ $=\binom{m}{2}\times\Big{(}1-\frac{1}{2^{t}}\Big{)}$
which underlies the known cut-off limit for markov chains  [Diaconis, 1996, Bayer and Diaconis, 1992]. As consequence, under random backgrounds, effect generalizability requires $2^{t}$ effect samples, according to Definition.5, in a fixed or fast $U-X$ system. Another trivial example has ordinal probability mass $p=\{1,0,0,0,...\}$ concentrated in a single gap, which, as expected, generates no inversions. This leads to the limit $|\Pi_{n}|=0$ which describe samples with no EV non-increasing sequences. An unobserved sample $U$ close to $p=\{1/q,1/q,1/q,1/q,...\}$ ($m\geq q$) is capable of asymptotically enumerating $|\Pi_{n}|=q!$ ordering of unobserved causes in least time $t$. The number of permutations, for a given $t$, is decreasing as we depart the second (random $U$) and approach the first (deterministic $p=\{1,0,0,0,...\}$). For intermediary and more typical scenarios, what matters is the particular distribution $p$ of confounders across a sample’s gaps.
Other than Eq.(11), there are many ways to show that any departure from uniformity will only decrease the number of EV-increasing sequences in the sample. Consider first the case of a stationary gap probability $p_{t}$. A new background is generated every $m/q$ time. Since $m$ is known and fixed, the $p_{t}$ where each and every factor effect is observed under the same background can be written as $p_{t}=(m/q,m/q,m/q,...)$ (the conditional uniform distribution) which can, in turn, be written as $p_{t}/m=(1/q,1/q,1/q,...)$. A simple approximation is $p/m=1/q+d_{j}$ where $\sum d_{j}=0$. Define then $s=\sum d^{2}$ as a ’phase’ between ${U,X}$ systems. This type of multinomial distribution approximation appear in the context of the birthday paradox  [Nunnikhoven, 1992, Gail et al., 1979] and reflects the recursive solution,
$S_{q,m}\Big{\{}\,p=[p_{1},...,p_{m-1}]\,\Big{\}}$ $=m!\sum_{X_{(1)}<X_{(2)}<...<X_{(m)}}p_{i}\times...\times p_{m}$ (12)
$=S_{q,m-1}\{[p_{1},...,p_{m-1}]\}+p_{m}S_{q-1,m-1}\{[p_{1},...,p_{m-1}]\}$ (13)
where the first equation summing over all subsets of $U$ such that $X_{(}1)<X_{(}2)<...<X_{(}m)$ and the second with initial conditions $S_{1,j}=\sum_{l=1}^{j}p_{i}$, and $S_{j,j}=p_{1}\times...\times p_{j}$ for $j=1,2,...,m$. Assuming a finite $q$, with each new sample ($m<q$), a further $p_{j}$ (rank) is observed, further approximating the (unknown) $p$ ($m\geq q$). The $s$ index is simply an indicator of changes in $U{-}X$ ’phase’ during this process, and will be useful when testing hypotheses about their relative changes.
For the case where all in-sample variables have the same ordinal frequencies - with their effect observations illustrated by the square diagram, Fig.1(b, lower-right) - $d$ describes departure from ideal $U{-}X$ effect background enumeration ($d$ is subquadratic and $q,m$ coprime), Definition.5. When $d\approx 0$, there is no ’phase’ between $X$ and $U$, and the sample treatment-background assignment mechanism is fixed. When $U$ and $p$ are fixed and $d>0$, effect backgrounds can be enumerated, due to its additive effect on $b(t)$, $\pi_{0}:b(t)=t+d\bmod q$ When $U$ is changing but $p$ is fixed, $d$, instead, has a multiplicative effect, $\pi_{0}:b(t)=d^{-t}\bmod q$, on enumeration and quickly becomes random333as result of the recursion $(d^{-1}(d^{-1}...(d^{-1}k(\bmod q-1))...)))$ for each gap position $k$. (an observation often employed by random number generators  [Knuth, 1997]). The relationship between $d$ and randomness is a deep, and somewhat unresolved, issue in mathematics  [Bayer and Diaconis, 1992, Diaconis and Fulman, 2023]. In the context of effect estimation and $U{-}X$ systems, this is a specially relevant case (i.e., of many time varying confounders), as it characterizes most real-world data and is where out-of-sample generalization is most useful. In this case, we want guarantees that any treatment-background assignment selected after $\pi_{0}$ could be chosen with the same probability (i.e., any of their orderings is as ’random’ as any other).
Definition 6 ($U{-}X$ Effect Background Randomization).
Under changing backgrounds $Z/X$ and fixed $p$, effect treatment-background enumeration can become asymptotically ergodic, and thus EV-increasing with high-likelihood; for an initial effect background $\pi_{0}$, the sample sizes required for this condition depend on the treatment-assignment distribution $p$, and the number of inversions from $\pi_{0}$ the distribution $p$ can generate, Eq.(11) (Sect.4.2 Effect Background Randomization).
This leads to the proposed solutions here for causal effect estimation with increasing generalizability,
Proposition 2.
Counterfactual background effect enumeration (Definition.6) is EV increasing in fast $U{-}X$ systems, and background randomization (Definition.5) is EV increasing in slow systems.
4 Related Work
Non-parametric tests (like permutation tests) can often be used only under assumptions of strict independence among factors and no unobserved factors. Consider, for example, a scenario where two sample units are observed (but not randomized), and we pair each unit receiving a first treatment $a$ with a different unit receiving $b$. To test the null hypothesis that the two treatments have no difference on any unit, we can permute treatment assignments within the pairs. The strategy relies on the assumption that the two units in the pair are surely exchangeable besides their treatment status. This assumption can be violated if the units are different in any unobserved way (even in high-dimensional samples it is hard for researchers to make the claim that there are no overlooked and relevant distinctions). Randomization plays a crucial role in alleviating the effects of unobserved confounding  [Zhang and Zhao, 2023, Kempthorne and Doerfler, 1969]. Randomized Experiments provide a solid foundation, and are still the golden rule in the Sciences, for making causal inferences about the effects of treatments on an outcome. Non-experimental samples can be seen as experimental samples where randomization did not occur perfectly. This leads to a scenario where treatment or control group are different, with units in one of them having factors that are predictive of (or are merely correlated with) the outcome, thus making treatment-control outcome differences invalid effect estimates. In these scenarios, standard non-parametric techniques becomes inapplicable. Despite these known shortcomings, non-parametric tests, specially permutation-based tests like the Shapley value  [Ribeiro, 2022a, Burkart and Huber, 2021], continue to be used in practice in non-i.i.d.and partially observed data, with little clarity over the resulting sample biases and generalizability.
The previous discussions illustrated the need to also permute (unobserved) effect backgrounds when estimating effects. According to our view, due to the inherent unobservability of those factors, the challenge for estimation then becomes how to establish in-sample conditions indicating when backgrounds (1) have not changed, or, (2) that they have changed enough such that any association between treatment and backgrounds of two effect observations or units nearly vanished. We suggested that a way to establish such conditions when $|\Pi_{n}^{*}(Z/\{a\})|\gg 0$ is to seek assurances of randomness from established theory, such as from Ergodic and Combinatorial Randomization theory, as further discussed below. In respect to the generalization of effects, another suggestion was that there are advantages in carrying randomization jointly across all $X$ factors.
Our main goal has been to formulate a quantifiable model for the generalizability of effect observations across populations. A lack of understanding about the issue creates artificial issues that surface once and again in non-parametric experimental studies, observational causal estimation, and black-box variable importance attribution. For example, Ribeiro et al.  [F. Ribeiro et al., 2022] showed that lack of sample variation explains one of the most heated methodological debates in Economics. Grimmer et al.  [Grimmer et al., 2020] mentions how the presence of ’Stan Lee’ in the credits of all Marvel franchise movies, as factor when estimating effects on movie success, leads celebrated algorithms like the Deconfounder  [Wang and Blei, 2020] to drastically overestimate the factor’s effect. More than any particular algorithm, what seems missing is a sound way to quantify the EV of effect estimates, and a better understanding of the issue altogether.
Concerns about robustness and predictive performance can, in part, be seen as the motivation for more recent views of causality based on model-performance invariance  [Peters et al., 2016, Magliacane et al., 2017, Buehlmann, 2020]. This is typically the invariance of a black-box model’s performance - and not of an effect, or, difference in outcomes, in a population, as taken here. The unit of analysis here, $\Delta y$, is much simpler (being an observation and not an estimate). Furthermore, on the present view, it is not the minimum variance of effects that matters, but the minimum variance of effects in face of the maximum variance of extraneous factors. The approach here addresses the latter conditions, and shed light on when successive observation of effects can lead to increases in out-of-sample performance. Finally, we believe the widespread success of the counterfactual definition is largely due to its simplicity, and simplicity has been a central guideline in the present work.
The issue of the Extenal Validity (EV) of effects has been addressed in  [Pearl and Bareinboim, 2014, 2011, Correa and Bareinboim, 2020]. Authors assume a known ’selection’ variable set $S$ that ’may represent all factors by which populations may differ or that may threaten the transport of conclusions between populations’ when training parametric models on non-experimental data. They then, assuming causal graphs for all other variables are known or inferred for both population, propose graphical conditions for EV based on Pearl’s graphical calculus. Assuming all this prior knowledge is known, the proposed conditions may give assurances on the transportability of ’observational findings’. Often those differences are not known. We take a more practical starting point, where every variable is potentially a relevant difference, no independence relations are assumed, and all results follow from in-sample effect differences and probabilities. We assume that subsets of ’population’ factors are observed ($X$) but another ($U$) is completely unobserved, as typical in supervised training tasks. We provide a quantifiable measure of EV (not a set of ideal graphical and independence conditions) for samples. In our view, the EV of effects is an uncertainty measure, Eq.(4), and not an isolated issue. At the same time, it is difficult to think of issues of external validity separate of issues of selection bias (which is the central issue in scientific research and Experimental Design)  [Cox, 2009, Montgomery, 2001]. The solution is formulated at the level of individual effect observations, non-parametrically, and in the counterfactual framework. We consider whether connections to Combinatorial Randomization might be helpful. The alternative combinatorial model proposed allow us to discuss critical issues - difficult to address, even if only theoretically, in the previous - like omitted variables, in-sample and out accuracy trade-offs, selection bias, prediction-interpretability tradeoffs, empirical results, and sample size requirements.
4.1 Randomization and Causal effects
We will use the notation often employed in Structurual Causal Models (SCM) to relate the previous concepts to current research. In SCM models, variables are typically specified by an unknown function of a set of endogeneous variables and an exogeneous error. We have, similarly, specified DGPs $Z=\{X,U\}$ with a collection of exogenous factors $U$, in addition to the observed $X$. Factors in $U$ are unobserved and not affected by $X$, but affect $X$. More specifically, a SCM specifies a collection of functions $f_{a}$ indexed by endogenous variables, $f=(f_{a}:a,b,c,...,[m])$. The SCM also calls for the specification of the relation between random variables $U$ and $X$. This is done through a set of functions $f=(f_{W},f_{a},f_{y})$, where $W$ are the other (non-treated) in-sample factors, $W=X-\{a\}$, and $y$ is the outcome of interest. The values of $W$, $A$, and $Y$ are then deterministically assigned by $U$, according to  [van der Laan and Rose, 2011]
$W=f_{W}(U_{W})$ (14)
$A=f_{a}(W,U_{a})$
$Y=f_{y}(W,A,U_{y})$
which defines a treatment-background assignment mechanism (without assumptions about $f_{W}$,$f_{A}$, and $f_{y}$).
Randomized controlled trials (RCTs) are one type of controlled experiment where subjects are randomized to receive a specific treatment. For example, if each subject was assigned to one of two levels of treatment based on the flip of a coin. Then the differences between the two groups would be solely due to treatment, as all other factors would be balanced (free os selection bias), up to random error. In an RCT the treatment assignment process, $P(A=1|W)$, is known (e.g., constant and $0.5$ as in a coin flip). This is very different from non-experimental data where treatment, or treatment exposure, is not assigned. In SCM notation, a RCT can be, nonetheless, written as in Eq.(15-16,18),
$W$ $=f_{W}(U_{W}),$ (15)
$A$ $=f_{a}(U_{a}),\;A\perp W$ (16)
[…]
$Y$ $=f_{y}(W,A,U_{y}).$ (18)
The only difference of Eq.(14) from Eq.(15-16,18) is the structural equation for $f_{a}$. The baseline variables $W$ play no role in the generation of $A$ in Eq.(16). These alternative assumptions can be briefly illustrated with causal graphs, Fig.2(c). Since $a$ is randomized, this implies that $U_{a}$ is independent of $U_{y}$ and $U_{W}$, and, in this case, no arrows connect $U_{a}$ to $U_{y}$ and $U_{a}$ to $U_{w}$.
4.2 Effect Background Randomization
Let $p=\{p_{1},p_{2},...,p_{m}\}$ be a sample gap background assignment. Consider a set of subsequent unobserved $U=Z\backslash X$ backgrounds $\Pi=\{\pi_{0},\pi_{1},\pi_{2},...\}$. Generalizability requires effects to be valid across any $\Pi$, irrespective of $p$. Eq.(8) stipulated two gap distributions that could allow us to estimate effects despite unobserved confounding. Imagine we keep $p$ constant from a time $t=0$ to $\delta t$ while allowing $\Pi_{n}(Z/X)$ to change. Can we say anything about whether the backgrounds $\pi_{0}$ and $\pi_{\delta t}$ are random (i.e., are as separated as in a uniform distribution), $\pi_{0}\perp\pi_{t}$? Fix $p$ and two backgrounds $\pi_{0},\pi_{1}\in\Pi_{n}(Z/X)$ separated by only one inversion. The likelihood of reaching $\pi_{1}$ from $\pi_{0}$ after one time step is high in this case. With increasing inversions in $\pi_{0}$, Eq.(11), there should be some period $\delta t$ where that likelihood becomes comparable to that of a uniform distribution - where $\binom{m}{2}$ inversions are generated at each time. Eq.(11) formulates that relationship. With an increasing $t$ (and $p$ with small $s$), this happens quickly, as its right-most term converges quickly to $1$. The choice of a background $\pi_{1}$ from $\pi_{0}$ is then as likely as any other background. Fig.2(b) shows this limit for the simulations in Sect.6 Experiments, which include correlated and non-stationary cases. The variation separation distance (y-axis) is a measure of randomness (i.e., formulated as a distance to an uniform distribution)  [Bayer and Diaconis, 1992, Diaconis, 1996]. The panel shows the distance between a fixed background $\pi_{0}$ and the subsequent in samples where $p$ is fixed and $s<0.25$ (these are exhaustively enumerated using the tests and procedures described next, distances where normalized to their per-factor range for display). The upper panel shows the same for other samples $s>0.75$. These combinatorial considerations suggest limits where some guarantees on effect background randomization can be made. Notice that $\binom{m}{2}$ in Eq.(11) is the number of effects observation in a single square. We will therefore use, more specifically, the number of inversions in squares to define testable conditions for background randomization in the next section.
[…]
The previous discussion indicates that, with random assignment of backgrounds to fixed treatments, as the sample size ($n$) approaches infinity, by the law of large numbers, the average characteristics of the treatment and control groups, $P(Z-\{a\}<a)$ and $P(Z-\{a\}>a$, converge to population averages. For a single factor $a$, we visualized these two groups as the upper and lower triangles of sample squares (i.e., in subsamples fulfilling the previous conditions). This can be written as
$\lim_{{n\to\infty}}P(Z-\{a\}<a)=P(Z-\{a\}>a).$ (19)
With convergence, the potential bias introduced by correlations in unobserved or uncontrolled factors ($U$) is mitigated. On average, any unobserved or uncontrolled factors that could potentially bias effect estimates will be equally distributed between the two effect observation orders, as required by Principle.1. In a square, this is true for all in-sample variables simultaneously ${\rm I\kern-1.99997ptE}\{\textrm{Var}[\,\Delta y(a)\,]\}\approx 0$. Notice that according to Eq.(19) this also implies that variables are expected to be confounded in equal amounts in this limit.
Proposition 3.
With random unobserved effect backgrounds and sufficiently large sample sizes $n$, subsample variables $a\in X$ converge asymptotically to true population treatment effects, despite $U{-}X$ correlations.
Availability of effect observations with these characteristics (and thus which strategy to use) is to be determined by statistical hypotheses tests. Hypotheses tests for these three conditions are similar but increasingly restrictive (i.e., requires larger samples).
4.3 Square Enumeration and Gap Inversion Tests
All three previous scenarios, and Principle.1, are related to permutation inversions. They were that $\pi_{0}$ changes, $\pi_{0}$ changes while $p$ stays constant for $\delta$ time, and $\pi_{0}$ changes while $p$ stays constant for $m\times\delta$ time. Inversions appear in common tests of i.i.d. sampling (’stationarity’) in ordinal data, such as in Kendall’s $\tau$ and Man-Whitney (when $X$ is binary) tests. Any pair of observations $(x_{i},y_{i})$ and $(x_{j},y_{j})$, where $i<j$, are said to be ’concordant’ if the sort order of $(x_{i},x_{j})$ and $(y_{i},y_{j})$ agrees: that is, if either both $x_{i}>x_{j}$ and $y_{i}>y_{j}$ holds or both $x_{i}<x_{j}$ and $y_{i}<y_{j}$; and be ’discordant’ otherwise. The number of discordant pairs between $x$ and $y$ is equal to the number of inversions that permutes the $y$-sequence into the same order of $x$. The statistic can also be seen as a generalised measure of a classifier’s separation power for more than two classes  [Hand and Till, 2001].
We formalize the previous conditions for non-experimental samples as any of their subsamples $C$, $C\subset X$, passing the following statistical hypothesis tests over inversions in sample ordinal and gap statistics,
[…]
Principle.1 is also related to inversions and says that confounders, causes and spurious factors have distinct effect variances under increasing inversions (e.g., the number of times confounder $b$ appears before its root cause). According to the principle, the observation of one inversion allows for the separation of a single cause-confounder pair (Appendix.A,B). We can generalize the previous conditions to non-experimental samples which are not sequentially sampled by enumerating squares for a specific sample unit $x_{0}$ using inter-unit combinatorial properties. In fully observed samples, $U=\emptyset$, square enumeration ensures the applicability of Principle.1. In samples with unobserved factors, $U\neq\emptyset$, enumeration with the additional conditions over gap inversions in Eq.(20) allows the same. Much like we can observe single factor differences in samples, and their effects, according to square enumeration, we can observe factor permutations and their effects, by combining single differences and their effects carefully. The panel in Fig.2(d) illustrates this process for two factors ($m=2$). It illustrates the differences (circles) and intersections (gray), from $x_{0}$, that characterize each cell in a square of size $2$. We say that by observing differences $a$, $b$ and $ab$ from $x_{0}$, we ’observed’ a permutation for these factors. That’s because these differences - $(x_{0}{-}x_{1})=\{a\}$, $(x_{0}{-}x_{2})=\{b\}$ and $(x_{0}{-}x_{3})=\{ab\}$ allow us to recover the effects of permuting factors $\{a,b\}$ (conditional on $x_{0}$), which are not directly observed in the sample. Notice that not all sample pairs with differences with $\{a\}$ and $\{ab\}$ would belong to the second square column, as there are also conditions on their intersection, $(x_{0}{-}x_{1})\cap(x_{0}{-}x_{3})=\{b\}$. It is the fulfillment of both combinatorial conditions that allow us to take them as (conditional) effect observations. The square full combinatorial structure generalizes this to $m>2$ and uses the $2^{m}$ $m$-way differences potentially available in samples. Further combinatorial, computational and statistical aspects of this enumeration process are described in Appendix.A,B,I.
Fig.1(a) summarizes this solution for effect estimation, where square enumeration ensures the inversions required for Principle.1 in in-sample variables, and tests over their their gaps, Eq.(20), in out-of-sample. For sets of units (and thus mutual effect observations) where these combinatorial and gap conditions hold simultaneously, variance in effects can become asymptotically indicative of the true status of factors as causes, Principle.1, Eq.(3) of their effects, and Eq.(4) of their effects generalizability. Variance becomes, in such conditions, a measurement of ’noise’ for individual effect observations that also takes causal principles into consideration  [F. Ribeiro et al., 2022].
4.4 Non-parametric Effect Estimation
The way both treatment assigments $A$, Eq.(18), and effects $Y$ are calculated, Eq.(18), in the present approach is related to other non-parametric causal effect estimators. Target learning is a semi-parametric approach that starts by specifying the mechanism by which individuals were assigned to different treatment levels. It does not assume random assignments. Specifically, taking the effect of a factor $a$ as the target parameter, it’s two main components are a treatment assignment mechanism $Q_{0}$, and an empirical distribution $P_{w}$. The first defines the probability distribution of receiving a particular treatment (e.g., the conditional probability of treatment given observed factors) and the second is the observed distribution of $W=X-\{a\}$ considering only those individuals who received the specific treatment (i.e., the distribution of the observed data for individuals who were treated according to a particular assignment mechanism). The targeted learning estimator $\hat{\Delta y(a)}_{n}$ is obtained by solving the following optimization problem:
$\hat{\Delta y(a)}_{n}=\arg\min_{\theta}\left\{\frac{1}{m-1}\sum_{w\in X-\{a\}}L\left(\Delta y(a),U_{0},P_{w}(a)\right)\right\}$
Here, $L$ is the loss function over model-specific hyperparameters $\theta$ that quantifies the discrepancy between the true parameter $\Delta y(a)$, the data-generating distribution $Q_{0}$, and the empirical distribution $P_{w}$ induced by a specific treatment assignment. The optimization is typically performed using ensemble or black-box machine learning techniques. The optimization thus aims to find hyperparameter $\theta$ that minimizes the loss across all treatment assignment mechanisms, resulting in an estimator that is robust to different ways treatments may have been assigned in the non-experimental sample. The approach here has the same goal, but relies on a theoretical connection between treatment-background assignment mechanisms, in one hand, and their relation to generalizability and confoudness of effects on the other. The theory provides an answer to when specific effect observations can be said to have the later properties, conditional on a fixed mechanism (as identified by sample gap statistics). Our main goal is to demonstrate that under some mechanisms input samples are capable of generating effects that are externally valid. The same is true of parametric methods that employ black-box techniques across time to generate treatment-control pairs for effect estimation [Calonico et al., 2019, Abadie, 2021].
In respect to outcome functions, Eq.(18), a general non-parametric strategy is to condition outcomes on all in-sample factors and take the difference in means for each stratum. This results in a treatment effect within each stratum of sample factors. One might now estimate the causal effect of treatment as the average over all strata of these strata-specific treatment effects. This approach is significantly more efficient than the unadjusted treatment effect estimators. That’s because the Law of Total Probabilities allows such methods to control (in-sample) confounding without the assumption that the effects are constant across strata (levels of confounders included in the model). If $U=\emptyset$, the g-formula  [Keil et al., 2014] implements this idea. For example, an outcome could be expressed as
$P(y=1)=\sum P(y=1|W=w)\times P(w)$ (21)
which indicates that we are summing over each possible value of $W=X-\{a\}$, and $P(W=w)$ is the probability that $W$ takes on the value $w$ in the reference population and only $a$ is left to vary. This adjusted estimator of the effect of $a$ is a nonparametric maximum likelihood estimator for $Y$ for a single treatment and covariate. Drawbacks are the requirement that $U=\emptyset$ and the loss in performance in strata with small number of sample units. In Eq.(3), effects are conditioned on all their backgrounds (where order is also taken to matter). The strategy in Eq.(21) is often combined with conditional independence tests to identify plausible in-sample confounders. While we presently favor simplicity to highlight theoretical considerations and unobserved factors, elements of these approaches (such as further in-sample conditional independence tests of the g-formula and the supervised plug-in flexibility of targeted learning) could be further combined with the proposed.
5 Sample Sizes
Non-experimental samples may contain subsamples that are simultaneously unbiased and predictive. Finding such subsamples shifts the requirement from one of carrying out an experimental intervention to one of accruing sufficient sample sizes across required combinatorial conditions. The expected sample size necessary to sample a square, when sample factors have different frequencies, is related in a simple way to its ordinal statistics. The asymptotics for recommended sample sizes, $\tilde{n}$, are, in particular,
$2^{m-1}P[\,X_{(m-1)}\,]^{-1},\quad\text{and,}\quad\Big{(}2^{m-1}P[\,X_{(m-1)}\,]^{-1}\Big{)}\times\phi,$ (22)
for the first and many squares scenarios, and where $\phi=1.6180...$ is the golden ratio (Appendix.H has a full proof). The number of squares in observational data are, this way, determined by its rare factors, which demand samples with increasing sizes to match the number of permutations in their balanced counterparts (i.e., where factors have equal frequencies).
Figure 3: (a) observed permutations and limits (m=10, balanced, log-scale); (b) square sample-unit histograms, increasing $n$.
6 Experiments
We consider simulations with increasing statistical complexity and in real-world applications. In both cases, we study three distinct prediction problems in held-out samples with $n$ observations. The tasks are to predict
[…]
These correspond to out-of-sample prediction of outcomes, prediction of intervention effects, and marginal factor effect estimation (evaluated against the ground-truth). The first is the central problem in supervised Machine Learning, the second is related to effect estimation with counterfactual and matching estimators [Morgan and Winship, 2007], and the third to both causal effect estimation [Pearl, 2000], and supervised explainability  [Burkart and Huber, 2021]. Effect estimation is, in the present perspective, an extreme omitted-variable problem. The problem gets easier with the decrease of unobserved confounders and the proposed framework can also be used to study the performance of generic samples and models.
[…]
    stationary and equiprobable factors with constant effects, (balanced)
    $p(a)=0.5,\>\Delta y(a)=1,\>\forall a\in X_{10}$ and $\forall u\in U_{3}$,
    non-stationary and factors with distinct probabilities and effects,  (unbalanced)
    $p(a)\sim\mathcal{U}([0,1]),\>\Delta y(a)\sim\mathcal{U}([0,1]),\>q\sim\mathcal{U}([3,20]),\>\forall a\in X_{10}$ and $z\in U_{q}$,
    non-stationary and correlated factors,  (correlated, ×$\times 3$)
    $p(a)\sim\mathcal{U}([0,1]),\>\Delta y(a)\sim\mathcal{U}([0,1]),\>q\sim\mathcal{U}([3,20]),\>\rho(a,b)=\mathcal{U}([0,\{0.1,0.25,0.5\}]),\>b=next(a),\>\forall a\in X_{10},$ and $\forall z\in U_{q}$,
    where $next(a)$ is $a$’s subsequent letter in lexicographic order,
For the last two cases (unbalanced and correlated), we also divide the sampling period into 10 time sub-periods, where at the beginning of each period the number of unobserved factors and parameters are re-sampled, while in-sample effects remain constant.
For $m$ factors and $d$ fixed-points, the number of permutations, derangements (permutations without overlaps), and partial permutations (permutations with $d$ positions fixed)  [Hanson et al., 1983] are, respectively,
$m!,\quad D_{m}=m!\sum_{d=0}^{m}\frac{(-1)^{d}}{d!},\quad\textrm{and},\quad D_{m}^{d}=\binom{m}{d}D_{m-d}.$ (23)
The last number indicates that, to form a partial permutation, we select $d$ unique factors to fix and be organized in $\binom{m}{d}$ ways, each with $D_{m-d}$ possible disjoint orderings of the non-selected factors. Fig.3(a) shows the number of observed permutations present on increasingly larger samples (balanced). It illustrates how sample sizes affect the number of permutations asymptotically enumerable in them. Each (green) curve corresponds to the number of permutations in samples of size $n=\{200,350,500,1000\}$ (blue curves show the number of squares). The lower panel shows combinatorial limits (red) for the number of partial permutations and squares according to Eq.(23), $m=10$. The figure suggests, more specifically, necessary sample sizes $n$ such that the number of permutations $\Pi_{n}(Z)$ coincide with the limit of all partial permutations555since $q\leq 20$ in these illustrations all permutations in a sample can be enumerated with the approach in Appendix.I, for each number $d$ of fixation-points, $d\leq m$. The relationship between the DGP and enumerable permutations or their speed of enumeration is, for example, formulated in Eq.(2). We expect these non-parametric counts to impact biases and predictive performance, Eq.(3), irrespective of the DGP. The same is shown as histograms of square member counts, Fig.3(b). At $n{=}200$, no differences of size larger than $4$ are observed. A full square of size $10$ is completely observed only at $n{=}1024$ (bottom-right) in the balanced case, and $n\approx 20K$ in the unbalanced case. This was anticipated by Eq.(22) and follow from simple sample ordinal statistics. The first sample size is $\tilde{n}=2^{9+1}=1024$. The latter sample has rare factors, ${\rm I\kern-1.99997ptE}[min_{a\in X}(p(a))]=0.025$, which leads to an approximate sample requirement of $\tilde{n}=2^{9}/0.025=20480$. Similar sample size recommendations will be indicated for each of the increasingly complex simulated and real-world cases considered below.
6.1 Outcome Prediction
Figure 4: supervised ACC vs. $n$ for the (a) unbalanced and (b) balanced cases; (c) ACC under different transversals; (d) a square transversal.
An understanding of the relationship among samples and estimators’ ACC and EV is necessary to devise accurate and maximally general models. To that end, we consider ACC (percentage of correctly classified cases in sample validation sections) as we increase $n$. For validation, we divide samples in two random sections: internal and external. The internal section is divided in typical training and validation subsections. We report accuracy of cross validation in the internal section (green) and external (red), with $4$ folds across cases. When ordered arbitrarily, or randomly, the internal-external division is inconsequential (as long as individual sections contain enough samples). We will define, however, alternative sample transversal orders, and consider how they impact learning performance. Starting with $n{=}1$, we train several algorithms with increasing $n$, and under different orders. With each unitary $n$ increase, an observation in the external section is transferred to the internal, and the problem of model and effect generalization becomes easier. Accordingly, a sample that can generalize its model and effects to external populations the earliest, for the same number of observations, can be said to have higher EV. We use three pre-specified XGBoost GBM (Gradient Boosting Machine) models, a grid of GLMs (Generalized Linear Model), a Random Forest (DRF), LASSO and Ridge regressions, five pre-specified GBMs, a near-default Deep Neural Net, an Extremely Randomized Forest (XRT), a random grid of XGBoost GBMs, a random grid of GBMs, and a random grid of Deep Neural Nets, as well as Stacked Ensembles with all previous models. Searched parameters are listed in Appendix.F.
We start with the case of increasing EV (ACC in the external section). We expect a relationship between the number of squares in samples and their EV. According to Proposition.1 and Eq.(9), the square transversal order with fastest increase in ACC corresponds to a leftmost column first, $X_{(1)}$, transversal. This transversal is illustrated in Fig.4(d). We will consider other transversal orders, and their effects in respect to EV and CF, below. Performance of algorithms in samples with increasing sizes, under random and square transversals, are shown in Fig.4(a,b) for, respectively, the unbalanced and balanced cases. ACC Curves for 1000 simulation runs are shown for all algorithms (best parameter set), but bold curves mark the leader666we enumerate squares over the set $X$ of factors without the outcome to transverse samples, but repeat it for $X\cup\{y\}$ in each training fold; this leads to increase in ACC, without using data (outcomes) unavailable in typical validation protocols. (best algorithm and parameter-set combination). Error bars are shown only for the leader. For the random transversal case, ACC (both external and internal) remain at constant levels. Since sampling is random, algorithms minimize in this case the empirical, expected error in sample populations - which can reflect both population frequencies and sample selection biases. For the square case, there is a steep (linear) increase in EV. This indicates that maximizing the number of squares in data can lead to increases in EV. Learning in this case no longer minimizes the expected risk, but the risk for the combinatorial set of all subpopulations. In these cases, there is little difference in performance among different algorithms, with low variance across runs (bars), suggesting analytic limits and explanations. Recommended sample sizes for the many-squares asymptotic case, Eq.(22), are marked with a dotted line across figures (mean over runs).
Figure 5: supervised ACC vs. $n$ for the (a) correlated, $\rho=\{0.1,0.25,0.5\}$, and (b) counterfactual prediction cases.
The figure also shows an alternative ’Travel-Salesman’ order (rightmost). It is obtained by solving a TSP [^7] Metric TSP  [Cormen, 2001](pg. 1029), in the worst case it generates solutions that are twice as long as the optimal tour (TSP calculation repeated in every simulation run). : sample units are cities, and their factor difference counts are distances. This order shares an important characteristic with square transversals: differences between units are kept small in this sampling order. The orders differ, however, in an important way: differences in the TSP are arbitrary and non-cyclic (following empirical sample frequencies). In this case, EV is steeply reducing, in contrast to square enumerations, Fig.4(a,b). This is not due to algorithms becoming sensitive to sample noise - as typical in overfitting. It portrays the expected behavior from algorithms when supplied with an increasingly complex population, generating models from specialized to general. The algorithmic stack includes regularized solutions (LASSO and Ridge regressions) and Target Learning (Sect. 4.4 Non-parametric Effect Estimation, Super-learner training, Appendix.F).
6.2 Counterfactual and Correlated Outcome Prediction
Patterns in Fig.4(a,b) repeat across all considered generative models, but we observe increasingly larger ACC gains proceeding from the balanced to the correlated case, Fig.5(a) [^8] TSP order (rightmost): only the best mean performance run is shown.. Current supervised solutions are expected to perform well in samples whose variables are already independent and unbiased. Fig.5(a) shows increasing EV gains under increasing maximum correlations $\rho=\{0.1,0.25,0.5\}$ (correlated). Notice that, also as expected, ACC increases are linear in all cases (under square orders), Eq.(9), and that asymptotic sample bounds, Eq.(22), also hold across correlated cases. Finally, this illustrates how square enumeration allows correlated data to be ’put into play’ for contemporary predictive solutions.
While the concepts above can give researchers larger control over the generalizability of learning solutions, we started with the goal of studying the EV of counterfactual predictions. This case is show in Fig.5(b). We take this to be the prediction of effect differences, $\Delta y_{ij}$, from covariate contrasts, $x_{i}-x_{j}$, for all validation unit pairs. The figure illustrates that everyday algorithms are clearly ineffective in this case. The intuition, and why the problem is rarely framed this way, was articulated above: using factor differences leads to samples with increased pairwise overlaps and decreasing permutations, decreasing the capability of algorithms to generalize effectively. The figure shows, however, that models can generalize effectively if these overlaps are considered carefully, as in square enumerations. This echoes the increased performance observed over correlated data, and reveals a relationship between these two problems.
Figure 6: causal and importance estimation Mean-Squared Error (MSE) vs. $n$, $\rho=\{0.1,0.25,0.5\}$, under (a) random order, (b) square-column order; (c) MSE under distinct transversals; (d) a square-column transversal.
6.3 Causal Effect Estimation and Black-box Explanations
Supervised solutions are highly tuned to generalizability. We now consider implications to causal techniques, which, instead, emphasize biasedness and estimation under non-i.i.d. conditions. Fig.4(d) and Fig.6(d) illustrate two distinct ways to transverse enumerated squares. Each individual square column corresponds to sets of effect observations with the same ordinal statistics and thus similar confounding level and variance, Proposition.1. Distinct square transversals reproduce EV increasing (Definition.2) and CF non-increasing (Definition.3) sequences. Imagine we place squares side-by-side, and that each square cell contains a sample unit. We transverse squares from left-to-right and top-to-bottom. The right-ward sequence of columns is, however, generated in one of two ways: across squares (a square transversal, like in the previous sections) and within squares (a square-column transversal). Using input from these two transversals should make estimators behave asymmetrically in respect to EV and CF. A square transversal corresponds to an order with increasing $EV$ and minimal $CF$ increase. A square-column transversal corresponds to an order with decreasing $CF$ and minimal $EV$ increase. The enumeration production rule for a column $d_{t}$, at time $t$, for each case is also shown in the figures, $d_{1}=1$. Fig.4(c) shows how ACC changes for supervised systems under random, square, and square-column transversals. Fig.6(c) shows how ACC changes, instead, for contemporary causal effect estimators and explaining systems (discussed below). Accuracy is, in this case, sum of squared differences between estimated and ground truth single-variable effects (for all variables). For a set of enumerated squares, for references $x_{0},x_{1},...,x_{t}$ at time $t$, accuracy is calculated for populations $x_{0},x_{1},...,x_{t}$. Lines correspond to mean ACC (and errors) of the best performing algorithm and parameter set (leader) across 1000 runs. Both transversals (square and square-column) supply systems with equally-represented populations but their performance is symmetric in respect to ACC. The random sample ordering strikes a balance between these extremes.
Figure 7: pairwise empirical errors, $(\Delta_{ij}y-\Delta_{uv}y)^{2}$, (y-axis) with increasing square sizes $m^{\prime}$ (color) and columns $d$ (x-axis), $d\leq m$, (a-d) balanced and unbalanced cases, (e-f) non-stationary case, (g-l) correlated cases (log-scale); (m) F-tests for balanced and unbalanced cases, and (n) for other cases (grey show results for subsamples passing tests in Eq.(20), omitted-correlated cases; (o) optimal sample split sizes; linear regressions, and confidence intervals (dashed ribbon), of effect error estimates for (p) omitted and (q) omitted-correlated cases.
Fig.6(a) shows performance of 3 widely used causal effect estimators and the IME explainer  [Lundberg and Lee, 2017, Lundberg et al., 2019] with random (top) and square-column (bottom) orders for three levels of sample correlation (correlated, square-column transversal)999with $n<100$, omitted datapoints are due to regression non-convergence.. The first estimator is the recent Deconfounder  [Wang and Blei, 2020](Sect. 3.1, linear Bayesian factor model fit with Variational Bayes, logistic outcomes, and Normal priors). The second is g-formula  [Chatton et al., 2020], an efficient solution popular in Epidemiology (Sect. 4.4 Non-parametric Effect Estimation). IME uses the Shapley value calculated from the output of all supervised methods in the last section. A propensity score matching estimator  [Rosenbaum and Rubin, 1983] is included for its popularity and baseline significance in counterfactual solutions. These algorithms (causal and explainer) make different assumptions. The first focuses on biasedness, the second combines (biased and often heuristic) predictive regressions for model selection. As expected, causal effect estimators perform well across correlated cases, compared to the explainer, which makes i.i.d. assumptions, Fig.6(a). Effect estimators loose accuracy with increasing correlation, however, as increasingly less data is uncorrelated in samples and used by these methods.
The approach here also establishes a connection to variance-based sensitivity analysis. Sensitivity analysis decompose systems’ error to estimate individual factor contributions. The Sobol index  [Sobol, 2001] is defined from the conditional variance of a factor, $\nicefrac{{\text{VAR}{\rm I\kern-1.39998ptE}[y|x_{i}]}}{{\text{VAR}(y)}}=\nicefrac{{\text{VAR}(\beta)}}{{\text{VAR}(y)}}$, where $\beta$ are effect estimates (often from Analysis of Variance, ANOVA). Most explaining and sensitivity analysis approaches start with a functional for outcomes, $y=f(a,b,c,...)$ where input factors $a,b,c,...$ are i.i.d. Black-box explainers subject boxes (algorithms) to a large number of variations, while sensitivity analysis decomposes their errors. The goal for both is to indicate the importance of individual input variables, especially in respect to sample predictive performance. The Shapley value comes from coalitional game theory and was devised axiomatically. Beyond its strategic origins, Shapley’s axioms are important requirements to decompose changes in any system  [de Boer and Rodrigues, 2020]. As it is often used with random sampling, it is closely related to bootstrap and permutation methods for black-box regressions.
Squares were used here as means to represent sets of effect observations for a unit $x_{0}$, their permutations in samples, and thus define non-parametric statistics for the effects of sample factors. Our initial definition of effects, Eq.(3), led to consequent definitions and interpretations for the EV, Definition.2, and CF, Definition.3, of counterfactual observations and samples. The unit of analysis here is $\Delta y$, an outcome difference observation, and not the outcomes of regressions. Permutations are not random sample intake orders, but ’observed’ permutations in samples that fulfilled a set of combinatorial and gap conditions, offering additional guarantees over effect estimates. Eq.(3) thus uses the complete set of $m$-way observed differences in a sample to permute observed counterfactual outcomes and derive effect estimates. The full set of variations (EV) ${\Pi}_{n}(Z\/{a/})$ is the full set of EV-increasing sequences in samples (or their ’external variations’) for sample populations, which is often limited by sample sizes $n$, or even be asymptotically unobservable for some treatment-background assignments.
Two fundamental problems with both the previous explaining and sensitivity analysis approaches are the need for parametric forms for outcomes, and the assumption of i.i.d. variables  [Aas et al., 2021]. The latter is also a problem for standard non-parametric tests and generic permutation-based tests  [Hoeffding, 1948, Yamato and Maesono, 1986]. Fig.6(a) shows that correlations introduce considerable biases into these methods’ estimates of variable importance. Fig.6(b) shows, however, that both methods perform well under square enumerations. Performance increases in both fronts (effect estimation and importance attribution) in this case. In fact, in small samples, the latter dominate recent methods designed specifically for causality and correlated data (upper-right panels). This illustrates that, when samples are limited and effect estimates not externally valid, solutions aimed at generalization (given their inputs are not biased) can offer gains in causal effect estimation. The Shapley (teal curve) and Sobol indices (black) largely coincide under these conditions, as variance becomes an unbiased indicator of importance and EV, Eq.(4). This illustrates that simple estimators (which do not involve over 20 supervised methods, repeated many times for explanations) can also become effective under square transversals. Notice that both the Deconfounder and g-formula are highly-tuned to the generative models used, as both rely on logistic regressions. This exemplifies the attraction of non-parametric insights, and that, even when models are correctly specified, there is room for gains with the proposed framework.
Altogether, these results suggest that sample size increases alone can not lead to increases in the generalizability of algorithms. The number of EV-increasing sequences in the sample play a key role. At the same time, the number of CF non-increasing play a role for causal effect estimation (or other tasks related to the estimation of individual factor effect). This observation is in direct contrast with, and suggests significant penalties for, the i.i.d. assumptions of black-box and bootstrap sampling approaches. These methods can observe gains and control relevant statistical requirements by sampling squares, instead of individual observations.
6.4 Square Gap and Column Statistics
The panels in Fig.7 (a-l) illustrate the proposed empirical error decomposition for counterfactual effect observations, Eq.(9), in enumerated sample squares. These diagrams depict the mean empirical error among pairs, $(\Delta_{ij}y-\Delta_{uv}y)^{2}$, (y-axis) in their $d$-th column (x-axis). These are observed errors among sample units’ outcomes (in the same square and column), and not effect estimates. The top row, Fig.7(a-d), shows errors in the previous cases (balanced, unbalanced, both transversals). Larger $d\leq m^{\prime}$, lead to pairs exposed to increasing unobserved confounders, and variance in effects, for a fixed factor and square reference unit $x_{0}$. Square and square-column transversals are shown. As expected, Eq.(9), increasing square sizes then lead to error decreases among its counterfactual observations at regular and linear rates (lines with different colors). These errors, $\textrm{VAR}_{(k)}$, are depicted by the approximately constant separation between lines. They offer effect estimates under increasing factor variation and EV, Eq.(4). Enumeration of multiple squares, for the same references, reduces errors for individual population members.
F-tests are tests to compare variance of groups and can be used to demonstrate the proposed (ex post) sample stratification strategy. F-tests follow straight-forwardly from the previous error decomposition, Eq.(9) (Appendix.G). Fig.7(m) shows, in particular, results of F-tests for variance of effects for units in cases (a-d) from the same populations (non-stratified data). It shows that F-tests lead to false-negative identification of same effect from variances (outside the statistic $5\%$ significance level) across all cases (horizontal bars are whisker box-plot ranges). The figure also shows the result of tests for subsamples that pass the tests in Eq.(20) (gray), which lead, instead, to true-posititive outcomes in the same simulated data. Fig.7(n) reproduces these results for the non-stationary case, where the correction is more significant. We demonstrate further uses of these estimates in Sect.6.6 Real-World Example. These results are possible due to the limit on effect background association illustrated in Fig.2(b). Fig.7(p-q) shows Bayesian ANOVA regressions, and their confidence intervals (dashed ribbons), for $\textrm{VAR}_{(k)}$ across 1000 runs of unbalanced and correlated cases. Different from the previous, simulations had $3$ fixed levels of unobserved confounders ($q=\{0,10,20\}$), the ground-truth variance in effect observations according to Eq.(9) is shown in black. ANOVA regressions over the full sample (without the proposed combinatorial and gap conditions) are non-convergent. These simple regressions and statistics can help researchers better understand the completeness, and possible generalizability, of their samples and effect estimates.
6.5 The 80/20 Sample Split
The most universally used rule-of-thumb in ML practice is the 80/20 ratio when splitting samples into train and test subsamples. Although it has remained at the theoretical margins of ML research, largely as application to the problem of overtraining in neural networks, a natural question is: why this proportion? There are two common answers. The first is heuristic, as a consequence of the Pareto Principle. The principle states that ’80% of effects come from 20% of causes’  [Chen et al., 1994]. Causes are used here vaguely, but the principle explains many natural and artificial phenomena. The second answer is more precise, formulated as a scaling-law for the ratio  [Guyon, 1997, Bahri et al., 2021], generalizing  [Amari et al., 1995, Kearns, 1995]. They find that ’the fraction of patterns reserved for the validation set should be inversely proportional to the square root of the number of free adjustable parameters’. In essence, the optimal split is therefore determined by the number of unique factors in a sample, and not its gross number of observations. Eq.(22)(multiple-squares) derived simple bounds, $\tilde{n}$, for sample sizes. The pareto principle and ratio scaling-law offer similar recommendations. Fig.7(o) shows mean difference between recommended sizes, $\tilde{n}$, and the optimal split, $n_{opt}$, across all previous simulation cases. In each case, a sample with $n^{\prime}>100$ observations is divided into 100 random and increasing subsamples. Each subsample adds $n^{\prime}/100$ new observations to the previous and is used as training for the previous supervised methods. The optimal split $n_{opt}$ is taken as the subsample size with an inflection point in the best performance (ACC) across all methods (all samples had ACC peaks, if non-unique the smallest size was taken as optimal). Error is thus $(\tilde{n}-n_{opt})/n^{\prime}$. Although an asymptotic approximation, relating to the asymptotic number of permutations in unbalanced samples, Eq.(22) describes well trade-offs in relation to the EV of training samples. Further empirical illustrations of these combinatorial tradeoffs can be found in  [Ribeiro, 2022a, b]
Figure 8: (a) ACC vs. $n$ COVID19 infection out-of-sample prediction in the UK Biobank (UKB); (b) Leave-one-out prediction MSE in the Center for Disease Control dataset (CDC)(upper maps), with added $\varepsilon_{sq}$ error component (lower maps); (c) infection error decomposition for the most-impacted US counties (log-scale).
6.6 Real-World Example
The COVID19 pandemic continues to threaten the lives and livelihoods of billions around the world. As of the break of the pandemic, there was a rush and expectations for Machine Learning (ML) solutions to help inform policy and individuals’ decisions  [Li et al., 2020, Jin et al., 2021, Verity et al., 2020]. In practice, few ML approaches were truly useful, with SIR semi-deterministic models  [Cramer et al., 2021], or their specializations, often favored. This is, in part, due to the few guarantees current supervised solutions can offer in respect to sample biases and heterogeneity, especially when samples are limited (which coincide with when predictions matter the most). SIR models offer predictions at high aggregate levels - most often country and city levels - taking all individuals therein to be the same. Some sources of heterogeneity have been identified, age being the most obvious. We discuss in particular, how to make predictions with moderate and severe number of unobservabed factors and small $n$. We consider two data sources. The first is the UK Biobank  [Bycroft et al., 2018] (UKB): a dataset with $\sim$500K UK citizens, 100K infected, 5K variables. It includes a wide variety of variables - not only from individuals’ electronic medical records, but also sociodemographic, economic, living, behavior and psychological annotations. The attraction is its completeness, high-dimensionality, and level (individual). We also consider a data dump from the American Center for Disease Control (CDC)  [CDC-Dataset-, 2021] with $\sim$24M cases but limited variables (age, geographic, race, ethnicity, and sex). The interest is as platform to discuss the severe unobserved factors case. Binary outcome is COVID19 infection in both cases. Both datasets include information up to April 1, 2021.
Fig.8(a) shows performance of the previous supervised solutions in the UK Biobank  [Bycroft et al., 2018]. These achieve out-of-sample ACC of $\sim$80with random sampling. We see, under square ordering, a pattern in EV similar to those in the previous tasks. Taking a fourth of the sample leads to 10% increase in ACC over the unseen data. As before, the TSP ordering leads to a sharp decrease in EV in the first-half section. At a fourth of the sample, we have larger confidence over the amount of population predictive coverage the trained models offer. The 10 variables with highest EV, according to Eq.(4), and their biobank codes (parenthesis) were: overcrowded-household (26432), [pop-density (20118), traffic-intensity (24011), time-to-services (26433)], [health-rank (2178), smoking (20116), expiratory-flow (3064), age (24)], job-physical (816), risk-taking (2040). The first is a deprivation index marking individuals in overcrowded households. This is in line with research showing that infection for families happen in a logical-OR fashion (i.e., members multiply their individual exposures to the disease)  [Rader et al., 2020, Emeruwa et al., 2020]. We demonstrated that the presence of many enumerable permutations in samples allow correlated factors, and confounder in particular, to be separated from causes by effect invariance, Principle.1. Variables in brackets have $\rho>0.1$, and are listed in order of effect invariance. There are many squares of size $10$, which indicates that, despite correlations, there are enough observations to also parse out correlated variables’ effects. Population density is favored over neighborhood traffic intensity, and general health over age, for example. Further description of these variables can be found on  [UKB-Showcase-, 2021] from their codes.
Most data used for prediction after the pandemic onset were, however, simpler, with only case counts and a few demographic variables. The COVID19 Case Surveillance database  [CDC-Dataset-, 2021] includes patient-level data reported by US states to the CDC. We use the subsample with simultaneously non-missing age, sex, race, ethnicity and location (county-level). This leads to 10 binary variables and $\sim$11M cases. We additionally generated a set of baseline controls from the corresponding variables in the 2018 American census (county level). Fig.8(c) shows $F$-test estimates for the UKB and CDC samples. Compared to the CDC dataset, the UKB has a comprehensive set of variables and $F$ under $10\%$ significance level. The estimation indicates that the CDC sample has very large numbers of omitted variables. Fig.8(c) also repeats the previous omitted variable plots for the CDC dataset and its 5 counties with highest case counts (top). As expected, $\textrm{VAR}_{(k)}$ ANOVA estimates, Eq.(LABEL:eq-stan), vary significantly across locations (upper-left, with county names) but these have little impact over observed variable pairwise errors (lower lines), as they affect observed square subpopulations uniformly and can be proxied-out, Eq.(9). The perspective suggests that, regardless of the CDC sample’s severe shortcomings, errors $\textrm{VAR}_{(k)}$ can be useful by indicating degree of ’noise’ and EV, Eq.(4) in effect observations. Fig.8(b)(top maps) shows Mean-Squared Error (MSE) of a Leave-One-Out task for the 1489 American counties in the CDC sample, using 3 of the best performers in the previous tasks. Case information from all other counties is used to predict a given location’s COVID19 incidence (addressing the question of which other locations’ infection information could be used to derive optimal estimates for a specific location). MSE is the mean squared difference between the predicted and (held-out) local incidence. Fig.8(b)(lower maps) repeats the task adding $\textrm{VAR}_{(k)}$ to the inputs of these state-of-the-art estimators, which leads to significant gains across locations and algorithms. The lower box-and-whisker plot summarizes results for the Gradient Boosting Machine  [Friedman et al., 2000] (left-most maps). This illustrates that even in clearly confounded samples, where it would be unadvised to assign relative importance to any of its factors, EV estimates can still carry useful information, as error estimates, for subsamples and locations. Other real-world examples are discussed in [Ribeiro, 2022a].
7 Discussion and Conclusion
The central methodological challenge in the Sciences, Policy-making and Design remains the evaluation of counterfactual statements (Did this treatment caused the result of interest? Did this policy?) The counterfactual definition of effects formulated sample properties necessary for effects to be free of selection biases, Eq.(1). The counterfactual definition of effects, Eq.(1), is based on the non-variation of effect backgrounds. The strategy in this work has been, in contrast, to focus on conditions where the amount of unobserved variation is guaranteed to be large (EV-increasing) and shared among factors (CF non-increasing). This alternative definition, as effects that remain invariant in the face of large external variation, can lead, instead, to generalizeable effect estimates. Under these conditions, causes can also be separated from confounders using effect variances, Principle.1. These considerations led to the notion of effect backgrounds, and the conditions under which they become sufficiently randomized. The perspective also allowed us to reconsider problems that have proven difficult for the pairwise counterfactual formulation, such as their sample size requirements and unobserved variable biases.
To that end, we formulated a new Combinatorial and Modular Arithmetic model to connect sample incompleteness to loss of estimation performance. We formulated sample properties necessary for effect External Validity, Eq.(3), and their required sample sizes. They were used to define EV and variance statistics for individual counterfactual effect observations, as well as for sets of counterfactual observations (samples). Several concepts, only loosely justified in mainstream ML, such the as the hyperbolic functions  [Ribeiro, 2022a, LeCun et al., ] of neural networks, transformation group invariance  [Bietti and Mairal, 2019, Cohen and Welling, 2016], and 80-20 train-test sample split appears naturally in this purely combinatorial perspective. Cyclic permutations, and treatment-background ’convolutions’, Eq.(10), suggest a bridge between combinatorial and continuous multi-scale or frequency-based representations. The latter consists of one of the theoretical underpinning of neural networks and several other methods. This suggests a possible direction for the long-sought connection between causality and mainstream ML techniques  [Scholkopf et al., 2021]. We focused on confounding, external validity, and selection bias. Other concepts from the causal inference literature could be similarly implemented.
A merit of the proposed approach is thus bringing causal and predictive concepts closer. We demonstrated generalizability bounds for popular supervised predictors and black-box explainers [Burkart and Huber, 2021]. Supervised prediction, black-box explanation and effect estimation approaches should state, side-by-side with their estimates, the conditions on which they are expected to hold (sample sizes, correlations, completeness, etc.). This is unknown for most current causal effect estimation solutions, IME explainers, and most out-of-sample prediction solutions. Until these are stated, it is difficult to, in practice, trust their outputs in real-life conditions. The approach allows us to model effects in both linear and non-linear data without the, rarely know and difficult to infer, parametric models for outcomes or strict independence among factors. The approach is related to the widely-employed Shapley value  [Burkart and Huber, 2021, Ribeiro, 2022a] and known U-Statistics  [Hoeffding, 1948, Yamato and Maesono, 1986], specialized for effect observations (not regressions), and generalized to non-i.i.d and partially observed data. We discussed out-of-sample, counterfactual, correlated and omitted variables prediction, as well as causal effect estimation, in simulations and an important real-world example. Correlated and counterfactual prediction observed, in particular, significant accuracy gains. A central result demonstrated here is that the EV of interventions are, to some extent, predictable from combinatorial properties of the populations they act upon. To a broad audience, Machine Learning methods are inherently limited due to their lack of generalizability and selection bias guarantees, when compared to, for example, experimental hypothesis testing. We believe the work could help build further connections between predictive and causal techniques.
Appendix A Cause-Confounder Separation from Effect Observations
Appendix B Separation and Effect Variance
Appendix C Expected Number of Draws for All Binary Outcomes
Appendix D Number of Squares and Derangements
Appendix E Alternating Binomial Coefficients Sum to Zero
Appendix F Supervised Algorithms and Parameters
Appendix G Bayesian Hierarchical Sample Errors Estimation
Appendix H Sample Sizes
H.1 Multiple Squares
Appendix I Enumeration